What would be the pattern and the next number in this sequence?:
1, 5, 10, 25, 50, ___

Answer by richwmiller(17219)

next is 85, 130,185,250
(script capital g)_n(a_n)(z) = (5 z (-2 z^2+z-1))/(z-1)^3+11,5,10,25,50
We start with two numbers 1 and 5
odds will be by 10 evens by 5
1 *10=10
5*5=25
then switch
multiply by 5 and 10
10*5=10
25*10=250
Next number is 250 according to that pattern